A bullet is fired in a horizontal direction with a muzzle velocity of 300m/s. In absence of air resistance, how far will it have dropped in traveling horizontal distance of 20m

1
Expert's answer
2022-04-19T10:00:34-0400

The equations of motion of a bullet

"x=v_0t\\
y=gt^2"

Hence

"y=gx^2/v_0^2=9.8*20^2/300^2\\
=0.044\:\rm m=4.4\: cm"